# 3cm3cm2.3cm2.3cm

### v1

@@ 1. Introduction
Now we describe our duality in more concrete terms. Let ${\bf w}=(w_{1},\ldots,w_{d^{\prime}})$ and ${\bf z}=(z_{1},\ldots,z_{{\boldsymbol{\ell}}})$ be sequences of complex numbers, and let ${\boldsymbol{\xi}}=(\xi_{1},\ldots,\xi_{{d^{\prime}}})$ and ${\boldsymbol{\gamma}}=(\gamma_{1},\ldots,\gamma_{{\boldsymbol{\ell}}})$ be sequences of positive integers, where ${d^{\prime}}$ and ${\boldsymbol{\ell}}$ are some positive integers satisfying ${d^{\prime}}\leq d$ and ${\boldsymbol{\ell}}\leq p+q+m+n$; see Section 4.1 for details. Let $t$ be an even variable and write $t_{a}=t-a$ for $a\in{\mathbb{C}}$. We consider the direct sums of Lie (super)algebras
@@
Let $\mathcal{A}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})$ denote the Gaudin algebra for ${\mathfrak{gl}}_{d}$ with singularities of orders $\gamma_{i}$ at $z_{i}$, $i=1,\ldots,{\boldsymbol{\ell}}$, and let $\mathcal{A}_{{p+m|q+n}}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})$ denote the Gaudin algebra for ${\mathfrak{gl}}_{p+m|q+n}$ with singularities of orders $\xi_{i}$ at $w_{i}$, $i=1,\ldots,{d^{\prime}}$; see Section 3.4, Section 3.5 and (4.15). The algebras $\mathcal{A}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})$ and $\mathcal{A}_{{p+m|q+n}}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})$ are commutative subalgebras of the universal enveloping algebras $U\big{(}{\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big{)}$ and $U\big{(}{\mathfrak{gl}}_{p+m|q+n}({{\bf w},{\boldsymbol{\xi}}})\big{)}$, respectively. Also, for $1\leq i\leq{\boldsymbol{\ell}}$, the singularity at $z_{i}$ is said to be regular if ${\boldsymbol{\gamma}}_{i}=1$ and irregular otherwise. The singularity type (either regular or irregular) of each $w_{j}$ is defined similarly.
@@
In this paper, we also study the classical versions of Gaudin models. Let $\mathcal{S}\big{(}{\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big{)}$ and $\mathcal{S}\big{(}{\mathfrak{gl}}_{p+m|q+n}({{\bf w},{\boldsymbol{\xi}}})\big{)}$ be the supersymmetric algebras of ${\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})$ and ${\mathfrak{gl}}_{p+m|q+n}({{\bf w},{\boldsymbol{\xi}}})$, respectively. They are also Poisson superalgebras and may be viewed as classical counterparts of $U\big{(}{\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big{)}$ and $U\big{(}{\mathfrak{gl}}_{p+m|q+n}({{\bf w},{\boldsymbol{\xi}}})\big{)}$, respectively. Let $\overline{\mathcal{A}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})\subseteq\mathcal{S}\big{(}{\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big{)}$ be the classical Gaudin algebra for ${\mathfrak{gl}}_{d}$ with singularities of orders $\gamma_{i}$ at $z_{i}$, $i=1,\ldots,{\boldsymbol{\ell}}$, and let $\overline{\mathcal{A}}_{{p+m|q+n}}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})\subseteq\mathcal{S}\big{(}{\mathfrak{gl}}_{p+m|q+n}({{\bf w},{\boldsymbol{\xi}}})\big{)}$ be the classical Gaudin algebra for ${\mathfrak{gl}}_{p+m|q+n}$ with singularities of orders $\xi_{i}$ at $w_{i}$, $i=1,\ldots,{d^{\prime}}$; see Section 5.2 and Section 5.3. The classical Gaudin algebras $\overline{\mathcal{A}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})$ and $\overline{\mathcal{A}}_{{p+m|q+n}}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})$ are Poisson-commutative.
@@ 2.1. The general linear Lie (super)algebra
[…]
@@
$[E^{i}_{j},E^{r}_{s}]=\delta_{j,r}E^{i}_{s}-(-1)^{(|i|+|j|)(|r|+|s|)}\delta_{i,s}E^{r}_{j}\qquad\mbox{for $i,j,r,s\in\mathbb{I}$.}$
@@ 2.2. Column determinants and Berezinians
Here $\mathfrak{S}_{k}$ denotes the symmetric group on $\{1,\ldots,k\}$ and $l(\sigma)$ denotes the length of $\sigma$. The row determinant of $A$ is defined to be
@@
$d_{i}(A)=\begin{vmatrix}a_{1,1}&\ldots&a_{1,i}\\ \ldots&\ldots&\ldots\\ a_{i,1}&\ldots&\framebox{$a_{i,i}$}\end{vmatrix},$
@@
Let $(s_{1},\ldots,s_{m+n})$ be a sequence of 0’s and 1’s such that exactly $m$ of the $s_{i}$’s are 0 and the others are 1. We call such a sequence a $0^{m}1^{n}$-sequence. Every $0^{m}1^{n}$-sequence can be written in the form $(0^{m_{1}},1^{n_{1}},\ldots,0^{m_{r}},1^{n_{r}}),$ where the sequence begins with $m_{1}$ copies of 00’s, followed by $n_{1}$ copies of $1$’s, and so on. The set of all $0^{m}1^{n}$-sequences is denoted by $\mathcal{S}_{{m|n}}$.
@@
$A=\begin{bmatrix}W&X\\ Y&Z\end{bmatrix},$ (2.2)
@@
By hypothesis, $A$ is a matrix of type $(0^{n})$ with commuting entries. By Proposition 2.5, $\det(A)=d_{1}(A)\ldots d_{n}(A)$. It follows that $[\det(A)]^{-1}=d_{1}(A)^{-1}\ldots d_{n}(A)^{-1}={\rm Ber}\,^{{\bf s}_{1}}(A)$ as the elements $d_{i}(A)$ commute. ∎
@@ 2.3. Pseudo-differential operators
Let $\mathcal{A}$ be an associative unital superalgebra over ${\mathbb{C}}$, and let $z$ and $w$ be commuting even variables. We denote by $\mathcal{A}[\kern-1.1pt[z]\kern-1.1pt]$ and $\mathcal{A}(\!(z)\!)$ the superalgebras of formal power series and Laurent series in $z$ with coefficients in $\mathcal{A}$, respectively. Let $\mathcal{A}(\!(z^{-1},w^{-1})\!)$ denote the superalgebra of all formal series of the form
$\sum_{j=-\infty}^{s}\sum_{i=-\infty}^{r}a_{ij}z^{i}w^{j},$
where $r,s\in{\mathbb{Z}}$ and $a_{ij}\in\mathcal{A}$.
We let $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ be the set of all formal series of the form
@@
We may endow $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ with a superalgebra structure using the rules:
$\partial_{z}\partial_{z}^{-1}=\partial_{z}^{-1}\partial_{z}=1,\quad\partial_{z}^{i}z^{j}=\sum_{k=0}^{\infty}\binom{i}{k}\binom{j}{k}k!\,z^{j-k}\,\partial_{z}^{i-k},\quad\mbox{for $i,j\in{\mathbb{Z}}$.}$ (2.3)
Here, e.g., ${\binom{i}{k}:=\frac{i(i-1)\ldots(i-k+1)}{k!}}$. Furthermore, let $\mathcal{A}(\!(\partial_{z}^{-1},z^{-1})\!)$ be the set of all formal series of the form
$\sum_{j=-\infty}^{s}\sum_{i=-\infty}^{r}a_{ij}\partial_{z}^{i}z^{j},\qquad r,s\in{\mathbb{Z}}\quad\text{and}\quad a_{ij}\in\mathcal{A}.$
Using the rules (2.3) but with $z$ and $\partial_{z}$ replaced with $\partial_{z}$ and $-z$, respectively, we may define a superalgebra structure on $\mathcal{A}(\!(\partial_{z}^{-1},z^{-1})\!)$. The superalgebras $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ and $\mathcal{A}(\!(\partial_{z}^{-1},z^{-1})\!)$ are called the superalgebras of pseudo-differential operators over $\mathcal{A}$.

### v2

# Dualities of Gaudin models with irregular singularities for general linear Lie (super)algebras

@@ 1. Introduction
[…]
@@
Let $\mathcal{A}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})$ denote the Gaudin algebra for ${\mathfrak{gl}}_{d}$ with singularities of orders $\gamma_{i}$ at $z_{i}$, $i=1,\ldots,{\boldsymbol{\ell}}$, relative to $\bf w$} and $\boldsymbol{\xi}$}, and let $\mathcal{A}_{{p+m|q+n}}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})$ denote the Gaudin algebra for ${\mathfrak{gl}}_{p+m|q+n}$ with singularities of orders $\xi_{i}$ at $w_{i}$, $i=1,\ldots,{d^{\prime}}$, relative to $\bf z$} and $\boldsymbol{\gamma}$}; see Section 3.4, Section 3.5 and (4.15). The algebras $\mathcal{A}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})$ and $\mathcal{A}_{{p+m|q+n}}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})$ are commutative subalgebras of the universal enveloping algebras $U\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)$ and $U\big({\mathfrak{gl}}_{p+m|q+n}({{\bf w},{\boldsymbol{\xi}}})\big)$, respectively. Also, for $1\leq i\leq{\boldsymbol{\ell}}$, the singularity at $z_{i}$ is said to be regular if ${\boldsymbol{\gamma}}_{i}=1$ and irregular otherwise. The singularity type (either regular or irregular) of each $w_{j}$ is defined similarly.
@@
In this paper, we also study the classical versions of Gaudin models. Let $\mathcal{S}\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)$ and $\mathcal{S}\big({\mathfrak{gl}}_{p+m|q+n}({{\bf w},{\boldsymbol{\xi}}})\big)$ be the supersymmetric algebras of ${\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})$ and ${\mathfrak{gl}}_{p+m|q+n}({{\bf w},{\boldsymbol{\xi}}})$, respectively. They are also Poisson superalgebras and may be viewed as classical counterparts of $U\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)$ and $U\big({\mathfrak{gl}}_{p+m|q+n}({{\bf w},{\boldsymbol{\xi}}})\big)$, respectively. Let $\overline{\mathcal{A}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})\subseteq\mathcal{S}\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)$ be the classical Gaudin algebra for ${\mathfrak{gl}}_{d}$ with singularities of orders $\gamma_{i}$ at $z_{i}$, $i=1,\ldots,{\boldsymbol{\ell}}$, relative to $\bf w$} and $\boldsymbol{\xi}$}, and let $\overline{\mathcal{A}}_{{p+m|q+n}}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})\subseteq\mathcal{S}\big({\mathfrak{gl}}_{p+m|q+n}({{\bf w},{\boldsymbol{\xi}}})\big)$ be the classical Gaudin algebra for ${\mathfrak{gl}}_{p+m|q+n}$ with singularities of orders $\xi_{i}$ at $w_{i}$, $i=1,\ldots,{d^{\prime}}$, relative to $\bf z$} and $\boldsymbol{\gamma}$}; see Section 5.2 and Section 5.3. The classical Gaudin algebras $\overline{\mathcal{A}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})$ and $\overline{\mathcal{A}}_{{p+m|q+n}}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})$ are Poisson-commutative.
@@ 2.1. The general linear Lie (super)algebra
Let $\{e_{i}\,|\,i\in\mathbb{I}\}$ be a basis for the superspace ${\mathbb{C}}^{p|q}\oplus{\mathbb{C}}^{m|n}$ such that $\{e_{i}\,|\,1\leq i\leq p+q\}$ and $\{e_{p+q+i}\,|\,1\leq i\leq m+n\}$ are respectively the standard homogeneous bases for ${\mathbb{C}}^{p|q}$ and ${\mathbb{C}}^{m|n}$. In other words, the parity of $e_{i}$ is given by $|e_{i}|=|i|$ for $i\in\mathbb{I}$.
@@
$E^{i}_{j},E^{k}_{l}]=\delta_{j,k}E^{i}_{l}-(-1)^{(|i|+|j|)(|k|+|l|)}\delta_{i,l}E^{k}_{j}\qquad\mbox{for $i,j,k,l\in\mathbb{I}$.}$
@@ 2.2. Column determinants and Berezinians
Here $\mathfrak{S}_{k}$ denotes the symmetric group on $\{1,\ldots,k\}$, and $l(\sigma)$ denotes the length of $\sigma$. The row determinant of $A$ is defined to be
@@
$d_{i}(A)=\begin{vmatrix}a_{1,1}&\ldots&a_{1,i}\\ \ldots&\ldots&\ldots\\ a_{i,1}&\ldots&\framebox{$a_{i,i}$}\end{vmatrix}$,\
@@
Let $(s_{1},\ldots,s_{m+n})$ be a sequence of 0’s and 1’s such that exactly $m$ of the $s_{i}$’s are 0 and the others are 1. We call such a sequence a $0^{m}1^{n}$-sequence. Every $0^{m}1^{n}$-sequence can be written in the form $(0^{m_{1}},1^{n_{1}},\ldots,0^{m_{r}},1^{n_{r}}),$ where the sequence begins with $m_{1}$ copies of 0’s, followed by $n_{1}$ copies of $1$’s, and so on. The set of all $0^{m}1^{n}$-sequences is denoted by $\mathcal{S}_{{m|n}}$.
@@
(2.2) $A=\begin{bmatrix}W&X\\ Y&Z\end{bmatrix}$,
@@
By hypothesis, $A$ is a matrix of type $(0^{n})$ with commuting entries. By Proposition 2.5, $\det(A)=d_{1}(A)\ldots d_{n}(A)$. It follows that $\det(A)]^{-1}=d_{1}(A)^{-1}\ldots d_{n}(A)^{-1}={\rm Ber}\,^{{\bf s}_{1}}(A)$ as the elements $d_{i}(A)$ commute. ∎
@@ 2.3. Pseudo-differential operators
Let $\mathcal{A}$ be an associative unital superalgebra over $\mathbb{C}$}, and let $z$ be an even variable. We denote by $\mathcal{A}[\kern-1.1pt[z]\kern-1.1pt]$ (resp., $\mathcal{A}(\!(z)\!)$) the superalgebra of formal power series (resp., formal Laurent series) in $z$ with coefficients in $\mathcal{A}$.
The superalgebra $\mathcal{A}(\!(z^{-1})\!)$ is equipped with an even derivation $\partial_{z}$ defined by
$\partial_{z}\left(\sum_{i=-\infty}^{r}a_{i}z^{i}\right)=\sum_{i=-\infty}^{r}ia_{i}z^{i-1},\qquad r\in{\mathbb{Z}}\quad\text{and}\quad a_{i}\in\mathcal{A}.$
Let $\mathcal{A}(\!(z^{-1})\!)(\!(\partial_{z}^{-1})\!)$ be the superspace of formal Laurent series in $\partial_{z}^{-1}$ with coefficients in $\mathcal{A}(\!(z^{-1})\!)$. The multiplication on $\mathcal{A}(\!(z^{-1})\!)$ extends to a multiplication on $\mathcal{A}(\!(z^{-1})\!)(\!(\partial_{z}^{-1})\!)$ by the rules
(2.3) $\partial_{z}\partial_{z}^{-1}=\partial_{z}^{-1}\partial_{z}=1,\quad\partial_{z}^{i}\phi=\sum_{k=0}^{\infty}\binom{i}{k}\,\partial_{z}^{k}(\phi)\,\partial_{z}^{i-k},\quad\mbox{for $\phi\in\mathcal{A}(\!(z^{-1})\!)$ and $i\in{\mathbb{Z}}$.}$
Here ${\binom{i}{k}:=\frac{i(i-1)\ldots(i-k+1)}{k!}}$ for $i\in{\mathbb{Z}}$ and $k\in{\mathbb{Z}}_{+}$.
Lemma 2.7.
$\mathcal{A}(\!(z^{-1})\!)(\!(\partial_{z}^{-1})\!)$ is an associative unital superalgebra over $\mathbb{C}$}.
Proof.
We need only show that the associative law for multiplication holds in $\mathcal{A}(\!(z^{-1})\!)(\!(\partial_{z}^{-1})\!)$. To establish the law, it suffices to show that
(2.4) $z^{i_{1}}\partial_{z}^{j_{1}}\big(z^{i_{2}}\partial_{z}^{j_{2}}z^{i_{3}}\partial_{z}^{j_{3}}\big)=\big(z^{i_{1}}\partial_{z}^{j_{1}}z^{i_{2}}\partial_{z}^{j_{2}}\big)z^{i_{3}}\partial_{z}^{j_{3}}$
for $i_{r},j_{r}\in{\mathbb{Z}}$, $r=1,2,3$. By (2.3),
$z^{i_{1}}\partial_{z}^{j_{1}}\big(z^{i_{2}}\partial_{z}^{j_{2}}z^{i_{3}}\partial_{z}^{j_{3}}\big)=\sum_{k=0}^{\infty}a_{k}z^{i_{1}+i_{2}+i_{3}-k}\partial_{z}^{j_{1}+j_{2}+j_{3}-k},$
where
$a_{k}=\sum_{r=0}^{k}\binom{j_{1}}{r}\binom{i_{2}+i_{3}-k+r}{r}\binom{j_{2}}{k-r}\binom{i_{3}}{k-r}r!(k-r)!,$
and
$\big(z^{i_{1}}\partial_{z}^{j_{1}}z^{i_{2}}\partial_{z}^{j_{2}}\big)z^{i_{3}}\partial_{z}^{j_{3}}=\sum_{k=0}^{\infty}b_{k}z^{i_{1}+i_{2}+i_{3}-k}\partial_{z}^{j_{1}+j_{2}+j_{3}-k},$
where
$b_{k}=\sum_{r=0}^{k}\binom{j_{1}}{r}\binom{i_{2}}{r}\binom{j_{1}+j_{2}-r}{k-r}\binom{i_{3}}{k-r}r!(k-r)!.$
Using Vandermonde’s identity:
$\binom{i+j}{r}=\sum_{s=0}^{r}\binom{i}{s}\binom{j}{r-s},\qquad i,j\in{\mathbb{Z}}\,\,\text{ and }\,\,r\in{\mathbb{Z}}_{+},$
we have for $k\in{\mathbb{Z}}_{+}$,
$a_{k}$ $=$ $\sum_{r=0}^{k}\binom{j_{1}}{r}\binom{j_{2}}{k-r}\left(\sum_{s=0}^{r}\binom{i_{2}}{r-s}\binom{i_{3}-k+r}{s}\right)\binom{i_{3}}{k-r}r!(k-r)!$
$=$ $\sum_{r=0}^{k}\sum_{s=0}^{r}\binom{j_{1}}{r}\binom{j_{2}}{k-r}\binom{i_{2}}{r-s}\binom{i_{3}}{k-r+s}\frac{r!(k-r+s)!}{s!}$
and
$b_{k}$ $=$ $\sum_{r=0}^{k}\binom{i_{2}}{r}\binom{i_{3}}{k-r}\binom{j_{1}}{r}\left(\sum_{s=0}^{k-r}\binom{j_{1}-r}{s}\binom{j_{2}}{k-r-s}\right)r!(k-r)!$
$=$ $\sum_{r=0}^{k}\sum_{s=0}^{k-r}\binom{i_{2}}{r}\binom{i_{3}}{k-r}\binom{j_{1}}{r+s}\binom{j_{2}}{k-r-s}\frac{(r+s)!(k-r)!}{s!}$
$=$ $\sum_{\ell=0}^{k}\sum_{s=0}^{\ell}\binom{i_{2}}{\ell-s}\binom{i_{3}}{k-\ell+s}\binom{j_{1}}{\ell}\binom{j_{2}}{k-\ell}\frac{\ell!(k-\ell+s)!}{s!}.$
Thus, $a_{k}=b_{k}$ for $k\in{\mathbb{Z}}_{+}$, and (2.4) is proved. ∎
The elements of the superalgebra $\mathcal{A}(\!(z^{-1})\!)(\!(\partial_{z}^{-1})\!)$ are called pseudo-differential operators. Let $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ denote the subspace of $\mathcal{A}(\!(z^{-1})\!)(\!(\partial_{z}^{-1})\!)$ consisting of all formal series of the form
@@
It is clear that $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ is closed under multiplication, and so it is a subalgebra of $\mathcal{A}(\!(z^{-1})\!)(\!(\partial_{z}^{-1})\!)$. Applying the rules (2.3), we see that any expression of the form ${\sum_{j=-\infty}^{s}\sum_{i=-\infty}^{r}a_{ij}\partial_{z}^{i}z^{j}}$, where $r,s\in{\mathbb{Z}}$ and $a_{ij}\in\mathcal{A}$, is an element of $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$. In particular,
(2.5) $\sum_{k=0}^{\infty}(-1)^{k}\binom{i}{k}\binom{j}{k}k!\,\partial_{z}^{j-k}\,z^{i-k}=z^{i}\partial_{z}^{j},\qquad\mbox{$i,j\in{\mathbb{Z}}$.}$
Conversely, any element of $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ can be written as ${\sum_{j=-\infty}^{s}\sum_{i=-\infty}^{r}a_{ij}\partial_{z}^{i}z^{j}}$ for some $r,s\in{\mathbb{Z}}$ and $a_{ij}\in\mathcal{A}$ due to (2.5).
It is easy to see from (2.3) and (2.5) that the identity map on the superalgebra $\mathcal{A}$ extends to a superalgebra automorphism
(2.6) ${\boldsymbol{\omega}}:\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$
such that ${\boldsymbol{\omega}}(z)=\partial_{z}$ and ${\boldsymbol{\omega}}(\partial_{z})=-z$. Also, $\boldsymbol{\omega}$} has order 4.
Let $z$ and $w$ be commuting even variables. For later use, we let $\mathcal{A}(\!(z^{-1},w^{-1})\!)$ denote the superalgebra of all formal series of the form
$\sum_{j=-\infty}^{s}\sum_{i=-\infty}^{r}a_{ij}z^{i}w^{j},\qquad r,s\in{\mathbb{Z}}\quad\text{and}\quad a_{ij}\in\mathcal{A}.$